OpenStudy (anonymous):

does tan^2x+sec^2x=1 for all values of x

OpenStudy (anonymous):

\[\tan^2 x+ \sec^2 x =\frac{ \sin^2x }{ \cos^2 x}+\frac{ 1 }{ \cos^2 x} = \frac{ \sin^2 x+1 }{ \cos^2 x } \neq 1\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Happy to help

OpenStudy (anonymous):

No how can you be sure that another x value would make statement wrong ?

OpenStudy (anonymous):

You need to prove for alll x values not for a particular x value

OpenStudy (anonymous):

Oh ! I am sorry. I didn't really give attention to your solution.

OpenStudy (anonymous):

So my solution is wrong or uncomplete ? What do you think ?

OpenStudy (anonymous):

What if you cannot find a particular x value that makes the statement wrong even if the the statement not true for all x values?